# Multistatic Radar Performance in the Presence of Distributed Wireless Synchronization

### v2

@@ III-A4 Adding Gaussian Priors for Offsets
The prior information matrix adds information only for the offset parameters $(\delta t_{i},\delta f_{i})$ (delay and Doppler are deterministic) and is expressed as
$\mathbf{J}_{P}^{i}=\operatorname{diag}(0,0,\sigma_{t_{i}}^{-2},\sigma_{f_{i}}^{-2})=\begin{bmatrix}0&0\\ 0&\boldsymbol{\Lambda}^{i}\end{bmatrix},\,\boldsymbol{\Lambda}^{i}=\begin{bmatrix}\sigma_{t_{i}}^{-2}&0\\ 0&\sigma_{f_{i}}^{-2}\end{bmatrix}$. (10)
@@
$\mathbf{J}_{B}^{i}=\mathbf{J}_{D}^{i}+\mathbf{J}_{P}^{i}=\begin{bmatrix}\mathbf{A}^{i}&\mathbf{A}^{i}\\ \mathbf{A}^{i}&\mathbf{A}^{i}+\boldsymbol{\Lambda}^{i}\end{bmatrix}$. (11)
@@ III-A5 Equivalent BIM for $\begin{bmatrix}\tau_{i}&f_{i}\end{bmatrix}^{\top}$
Eliminating the nuisance offsets using the Schur complement yields $\mathbf{J}_{\text{eq}}^{i}=\mathbf{A}^{i}-\mathbf{A}^{i}(\mathbf{A}^{i}+\boldsymbol{\Lambda}^{i})^{-1}\mathbf{A}^{i}.$
@@ III-A6 Bayesian CRLB
The BCRLB for the parameters $\begin{bmatrix}\tau_{i}&f_{i}\end{bmatrix}^{\top}$ is obtained as the inverse of the equivalent BIM
$\operatorname{CRLB}(\begin{bmatrix}\tau_{i}&f_{i}\end{bmatrix}^{\top})=\left(\mathbf{J}_{\text{eq}}^{i}\right)^{-1}.$ (12)
IV CRLB for Position $\mathbf{x}$ and Velocity $\mathbf{v}$
We consider the leader as the monostatic transmitter–receiver radar (indexed as $i=0$) and $N$ receive-only followers ($i=1,\dots,N$). The target position and velocity are denoted by $\mathbf{p}\in\mathbb{R}^{n}$ and $\mathbf{v}\in\mathbb{R}^{n}$, respectively, where $n\in\{2,3\}$. The transmitter is located at $\mathbf{p}_{t}$, receiver $i$ is located at $\mathbf{p}_{r,i}$, and the monostatic radar satisfies $\mathbf{p}_{t}=\mathbf{p}_{r,0}$. The target-to-radar unit vectors and corresponding ranges are defined as
@@ IV CRLB for Position $\mathbf{x}$ and Velocity $\mathbf{v}$
$\mathbf{y}_{i}\triangleq\begin{bmatrix}T_{i}\\ F_{i}\end{bmatrix}=\underbrace{\begin{bmatrix}\tau_{i}(\mathbf{p})\\ f_{i}(\mathbf{p},\mathbf{v})\end{bmatrix}}_{\boldsymbol{\mu}_{i}(\boldsymbol{\psi},\boldsymbol{\eta})}+\underbrace{\begin{bmatrix}w_{i}\\ v_{i}\end{bmatrix}}_{\mathbf{n}_{i}},$ (15)
where $\boldsymbol{\psi}=[\mathbf{p}^{\top}\ \mathbf{v}^{\top}]^{\top}$ represents the parameters of interest. The terms $w_{i}$ and $v_{i}$ denote the TOA and FOA measurement errors, respectively. The noise covariance for $\mathbf{n}_{i}\sim\mathcal{CN}(0,\boldsymbol{\Sigma}_{i})$ is determined from (12), i.e., $\boldsymbol{\Sigma}_{i}=\operatorname{CRLB}([\tau_{i},f_{i}]),$ which accounts for the both measurements and residual frequency and time offsets.
@@
Stack the measurement and mean vectors as $\mathbf{y}=[\mathbf{y}_{0}^{\top},\dots,\mathbf{y}_{N}^{\top}]^{\top},$ and $\boldsymbol{\mu}=[\boldsymbol{\mu}_{0}^{\top},\dots,\boldsymbol{\mu}_{N}^{\top}]^{\top}$, and since the measurements at each radar are statistically independent, the overall noise covariance matrix becomes a block-diagonal matrix $\mathbf{R}=\mathrm{blkdiag}(\boldsymbol{\Sigma}_{0},\dots,\boldsymbol{\Sigma}_{N-1})$. The full FIM can then be written as the sum of the contributions from each radar
$\mathbf{J}(\boldsymbol{\psi})$ $=\left(\frac{\partial\boldsymbol{\mu}}{\partial\boldsymbol{\psi}}\right)^{\!\top}\mathbf{R}^{-1}\left(\frac{\partial\boldsymbol{\mu}}{\partial\boldsymbol{\psi}}\right)$
$=\sum_{i=0}^{N}\left(\frac{\partial\boldsymbol{\mu}_{i}}{\partial\boldsymbol{\psi}}\right)^{\!\top}\boldsymbol{\Sigma}_{i}^{-1}\left(\frac{\partial\boldsymbol{\mu}_{i}}{\partial\boldsymbol{\psi}}\right)\triangleq\sum_{i=0}^{N}\mathbf{J}^{i}.$ (18)
@@ IV-A Partial Derivatives of the Mean
For the $i$th radar, the derivative of the mean vector with respect to the parameter vector $\boldsymbol{\psi}$ is given by $\frac{\partial\boldsymbol{\mu}_{i}}{\partial\boldsymbol{\psi}}=\mathbf{G}_{i}$, where $\mathbf{G}_{i}=\begin{bmatrix}\dfrac{\partial\tau_{i}}{\partial\mathbf{p}}&\dfrac{\partial\tau_{i}}{\partial\mathbf{v}}\\ \dfrac{\partial f_{i}}{\partial\mathbf{p}}&\dfrac{\partial f_{i}}{\partial\mathbf{v}}\end{bmatrix}\in\mathbb{R}^{2\times 2n}.$The derivatives $\frac{\partial\boldsymbol{\mu}_{i}}{\partial\boldsymbol{\psi}}$ are derived using the definitions in Appendix Appendix: Vector Derivatives. For a general unit vector defined as $\hat{\mathbf{u}}(\mathbf{p})=\frac{\mathbf{p}-\mathbf{a}}{\|\mathbf{x}-\mathbf{a}\|}$ with range $r=\|\mathbf{p}-\mathbf{a}\|$, the following key derivative identities hold
@@
$\frac{\partial\tau_{i}}{\partial\mathbf{p}}$ $=\begin{cases}\dfrac{1}{c}\!\left(\hat{\mathbf{u}}_{t}+\hat{\mathbf{u}}_{r,i}\right)^{\!\top},&i\geq 1\\ \dfrac{2}{c}\,\hat{\mathbf{u}}_{r,0}^{\top},&i=0\end{cases}\!,\hskip 18.49988pt\frac{\partial\tau_{i}}{\partial\mathbf{v}}=\mathbf{0}^{\top}.$ (20)
@@
$\frac{\partial f_{i}}{\partial\mathbf{v}}$ $=\begin{cases}\dfrac{1}{\lambda}\!\left(\hat{\mathbf{u}}_{t}+\hat{\mathbf{u}}_{r,i}\right)^{\!\top},&i\geq 1\\ \dfrac{2}{\lambda}\,\hat{\mathbf{u}}_{r,0}^{\top},&i=0\end{cases}\!,$ (21)
@@
$\frac{\partial f_{i}}{\partial\mathbf{p}}$ $=\begin{cases}\dfrac{1}{\lambda}\,\mathbf{v}^{\!\top}\!\left[\dfrac{1}{r_{t}}\!\left(\mathbf{I}-\hat{\mathbf{u}}_{t}\hat{\mathbf{u}}_{t}^{\top}\right)+\dfrac{1}{r_{r,i}}\!\left(\mathbf{I}-\hat{\mathbf{u}}_{r,i}\hat{\mathbf{u}}_{r,i}^{\top}\right)\right],&i\geq 1\\ \dfrac{2}{\lambda}\,\mathbf{v}^{\!\top}\!\left[\dfrac{1}{r_{r,0}}\!\left(\mathbf{I}-\hat{\mathbf{u}}_{r,0}\hat{\mathbf{u}}_{r,0}^{\top}\right)\right],&i=0\end{cases}\!.$ (22)
Finally, the FIM contribution from the $i$th radar reduces to $\mathbf{J}^{i}=\mathbf{G}_{i}^{\top}\boldsymbol{\Sigma}_{i}^{-1}\mathbf{G}_{i},$ and the total effective FIM for the target position and velocity parameters is obtained by summing the individual contributions $\mathbf{J}(\boldsymbol{\psi})=\sum_{i=0}^{N}\mathbf{J}^{i},$ and the corresponding CRLB for $\begin{bmatrix}\mathbf{p}&\mathbf{v}\end{bmatrix}^{\top}$ is given by
$\operatorname{CRLB}(\begin{bmatrix}\mathbf{p}&\mathbf{v}\end{bmatrix}^{\top})=\left[\mathbf{J}(\boldsymbol{\psi})\right]^{-1}.$ (23)
@@ V-1 Frequency Synchronization
The leader broadcasts a TTW, where the tones are separated by a fixed reference frequency, to all followers [1, 2, 15, 16, 5]. Each follower extracts the clock signal using a dedicated circuit and uses it as a reference for clock synchronization. The TTW generated at the leader is also based on the same reference frequency. This ensures that the clocks and carriers at all radars tick at the same rate. However, clock offsets may still remain unaligned.
@@ VI Simulation Results
[…]
@@
Leader / Follower transmit power, $P^{L}_{\text{sync}}/P^{F}_{\text{sync}}$ 33/33 dBm
@@
For the leader–follower synchronization links, an Urban Macro (UMa) large-scale path-loss model with Rician fading is adopted [10], capturing both dominant LOS and diffuse NLOS components (exponential delay power profile). Depending on the node heights, the channels are categorized as Air-to-Ground (A2G) or Air-to-Air (A2A). As the radar altitude increases, the LOS component becomes stronger, leading to higher SNR and reduced multi-path effects. A Rician factor of 3 is used following [10], with both leader and followers placed at a height of 10 m (A2A configuration). For frequency synchronization, the leader broadcasts a TTW with 10 MHz tone separation. Each follower extracts a reference clock from the received waveform and uses it to drive its internal circuitry. The frequency difference between the extracted reference and the nominal 10 MHz tone spacing is estimated and multiplied by the PLL (Phase-locked loop) factor to obtain the CFO error. The PLL factor refers to the multiplication ratio that a PLL applies to its reference input frequency to generate the carrier. For time synchronization, followers initiate a bidirectional waveform exchange at a transmit power of $P_{\text{sync}}^{F}$, and the leader responds accordingly. Since clock alignment depends on delay estimation (independent of the TOA estimation used for target tracking), the clock estimation error is modeled as twice the delay estimation error, accounting for estimation at both ends. An LFM waveform is used for delay estimation [15]. The CRLB for the first-path delay estimate is
@@
where $\zeta_{f}^{2}$ is the mean-squared bandwidth, the (pre-processed) $\mathrm{SNR}_{\mathrm{sync}}=P_{\mathrm{sync}}/\sigma_{n}^{2}$ with transmit power $P_{\mathrm{sync}}$ and noise variance $\sigma_{n}^{2}$, and $|h^{1}_{\mathrm{sync}}|^{2}$ is the dominant-path channel gain. For an LFM pulse of bandwidth $B_{\mathrm{sync}}$, $\zeta_{f}^{2}=\frac{(\pi B_{\mathrm{sync}})^{2}}{3}.$ Thus, the standard deviations of the frequency ($\delta f$) and clock time ($\delta t$) errors are computed and used as priors in the FIM calculation for delay and Doppler estimation at each radar. A time–bandwidth product of $B_{\mathrm{sync}}T_{\mathrm{sync}}=1$ is assumed.
The standard deviations of the CFO errors and clock error are shown in the Fig. 3. The maximum Doppler shift in the synchronization links is considered to be 0.1 Hz, arising from environmental disturbances. The performance plots show the clock error in nanoseconds (ns), with higher errors observed for followers farther from the leader and lower errors for those closer to it, indicating that higher SNR yields smaller clock errors. Increasing the $B_{\mathrm{sync}}$ further reduces the clock error. The CFO errors are primarily influenced by Doppler shifts and diffuse multi-path components in the environment. As multi-path severity increases, the likelihood of one or both tones experiencing deep fades also rises. Lower delay spreads correspond to higher coherence bandwidths; hence, higher CFO errors are observed for delay spreads exceeding 100 ns.
@@
Fig. 4 compares the MSR performance under ideal conditions and with synchronization offsets introduced by the integrated synchronization technique. The PEB (position error bound) and VEB (velocity error bound) are obtained from (23), computed as the square root of the first three and the next three diagonal terms, respectively, with target locations and velocities uniformly drawn from the ranges in Table III. Performance improves with increasing $B_{\mathrm{sync}}$ used for delay estimation, while stronger diffuse multi-path (increasing delay spread $\tau_{rms}$) degrade accuracy. As synchronization-link performance improves, the MSR approaches the ideal case. The PEB is also compared with that of the monostatic radar [7]. The VEB remains unaffected by the synchronization bandwidth, as it is only influenced by the CFO errors resulting from frequency synchronization.

### v3

@@ III-A4 Adding Gaussian Priors for Offsets
The prior information matrix adds information only for the offset parameters $(\delta t_{i},\delta f_{i})$ (delay and Doppler are unknown deterministic) and is expressed as
$\mathbf{I}_{P}^{i}=\operatorname{diag}(0,0,\sigma_{t_{i}}^{-2},\sigma_{f_{i}}^{-2})=\begin{bmatrix}0&0\\ 0&\boldsymbol{\Lambda}^{i}\end{bmatrix},\,\boldsymbol{\Lambda}^{i}=\begin{bmatrix}\sigma_{t_{i}}^{-2}&0\\ 0&\sigma_{f_{i}}^{-2}\end{bmatrix}$. (10)
@@
$\mathbf{I}_{B}^{i}=\mathbf{I}_{D}^{i}+\mathbf{I}_{P}^{i}=\begin{bmatrix}\mathbf{A}^{i}&\mathbf{A}^{i}\\ \mathbf{A}^{i}&\mathbf{A}^{i}+\boldsymbol{\Lambda}^{i}\end{bmatrix}$. (11)
@@ III-A5 Equivalent BIM for $\begin{bmatrix}\tau_{i}&f_{i}\end{bmatrix}^{\top}$
Eliminating the nuisance parameters using the Schur complement yields $\mathbf{I}^{\text{eq}}_{i}=\mathbf{A}^{i}-\mathbf{A}^{i}(\mathbf{A}^{i}+\boldsymbol{\Lambda}^{i})^{-1}\mathbf{A}^{i}.$
@@ III-A6 Bayesian CRLB
The lower bound for estimation error for the parameters $\tau_{i}$ and $f_{i}$ is obtained as
$\operatorname{CRLB}(\begin{bmatrix}\tau_{i}&f_{i}\end{bmatrix}^{\top})=\left(\mathbf{I}^{\text{eq}}_{i}\right)^{-1}.$ (12)
IV CRLB for Position $\mathbf{p}$ and Velocity $\mathbf{v}$
We consider the leader as the monostatic transmitter–receiver radar (indexed as $i=0$) and $N$ receive-only followers ($i=1,\dots,N$). The target position and velocity are denoted by $\mathbf{p}\in\mathbb{R}^{n}$ and $\mathbf{v}\in\mathbb{R}^{n}$, respectively, where $n$ is 3. The transmitter is located at $\mathbf{p}_{t}$, receiver $i$ is located at $\mathbf{p}_{r,i}$, and the monostatic radar satisfies $\mathbf{p}_{t}=\mathbf{p}_{r,0}$. The target-to-radar unit vectors and corresponding ranges are defined as
@@ IV CRLB for Position $\mathbf{p}$ and Velocity $\mathbf{v}$
$\mathbf{x}_{i}\triangleq\begin{bmatrix}T_{i}\\ F_{i}\end{bmatrix}=\underbrace{\begin{bmatrix}\tau_{i}(\mathbf{p})\\ f_{i}(\mathbf{p},\mathbf{v})\end{bmatrix}}_{\boldsymbol{\mu}_{i}(\boldsymbol{\psi})}+\underbrace{\begin{bmatrix}w_{i}\\ v_{i}\end{bmatrix}}_{\mathbf{n}_{i}},$ (15)
where $\boldsymbol{\psi}=[\mathbf{p}^{\top}\ \mathbf{v}^{\top}]^{\top}$ represents the parameters of interest. The terms $w_{i}$ and $v_{i}$ denote the TOA and FOA measurement errors, respectively. The noise covariance for $\mathbf{n}_{i}\sim\mathcal{N}(0,\boldsymbol{\Sigma}_{i})$ is determined from (12), i.e., $\boldsymbol{\Sigma}_{i}=\operatorname{CRLB}([\tau_{i},f_{i}]),$ which accounts for the both measurements and residual frequency and time offsets.
@@
Stacking the measurement and mean vectors $\mathbf{x}=[\mathbf{x}_{0}^{\top},\dots,\mathbf{x}_{N}^{\top}]^{\top},$ and $\boldsymbol{\mu}=[\boldsymbol{\mu}_{0}^{\top},\dots,\boldsymbol{\mu}_{N}^{\top}]^{\top}$, and since the measurements at each radar are statistically independent, the overall noise covariance matrix becomes a block-diagonal matrix $\mathbf{R}=\mathrm{blkdiag}(\boldsymbol{\Sigma}_{0},\dots,\boldsymbol{\Sigma}_{N})$. The full FIM can then be written as the sum of the contributions from each radar
$\mathbf{I}(\boldsymbol{\psi})$ $=\frac{\partial\boldsymbol{\mu}}{\partial\boldsymbol{\psi}}\mathbf{R}^{-1}\frac{\partial\boldsymbol{\mu}}{\partial\boldsymbol{\psi}}^{\!\top}=\sum_{i=0}^{N}\frac{\partial\boldsymbol{\mu}_{i}}{\partial\boldsymbol{\psi}}\boldsymbol{\Sigma}_{i}^{-1}\frac{\partial\boldsymbol{\mu}_{i}}{\partial\boldsymbol{\psi}}^{\!\top}.$
@@ IV-A Partial Derivatives of the Mean
For the $i$th radar, the derivative of the mean vector with respect to the parameter vector $\boldsymbol{\psi}$ is given by $\frac{\partial\boldsymbol{\mu}_{i}}{\partial\boldsymbol{\psi}}=\mathbf{G}_{i}$, where $\mathbf{G}_{i}=\begin{bmatrix}\dfrac{\partial\tau_{i}}{\partial\mathbf{p}}&\dfrac{\partial f_{i}}{\partial\mathbf{p}}\\ \dfrac{\partial\tau_{i}}{\partial\mathbf{v}}&\dfrac{\partial f_{i}}{\partial\mathbf{v}}\end{bmatrix}\in\mathbb{R}^{2n\times 2}.$ The derivatives $\frac{\partial\boldsymbol{\mu}_{i}}{\partial\boldsymbol{\psi}}$ are derived using the definitions in Appendix Appendix: Vector Derivatives. For a general unit vector defined as $\hat{\mathbf{u}}(\mathbf{p})=\frac{\mathbf{p}-\mathbf{a}}{\|\mathbf{x}-\mathbf{a}\|}$ with range $r=\|\mathbf{p}-\mathbf{a}\|$, the following key derivative identities hold
@@
$\frac{\partial\tau_{i}}{\partial\mathbf{p}}$ $=\begin{cases}\dfrac{1}{c}\!\left(\hat{\mathbf{u}}_{t}+\hat{\mathbf{u}}_{r,i}\right)^{\!\top},&i\geq 1\\ \dfrac{2}{c}\,\hat{\mathbf{u}}_{r,0}^{\top},&i=0\end{cases}\!,\hskip 18.49988pt\frac{\partial\tau_{i}}{\partial\mathbf{v}}=\mathbf{0}^{\top}.$
@@
$\frac{\partial f_{i}}{\partial\mathbf{v}}$ $=\begin{cases}\dfrac{1}{\lambda}\!\left(\hat{\mathbf{u}}_{t}+\hat{\mathbf{u}}_{r,i}\right)^{\!\top},&i\geq 1\\ \dfrac{2}{\lambda}\,\hat{\mathbf{u}}_{r,0}^{\top},&i=0\end{cases}\!,$
@@
$\frac{\partial f_{i}}{\partial\mathbf{p}}$ $=\begin{cases}\dfrac{1}{\lambda}\,\mathbf{v}^{\!\top}\!\left[\dfrac{1}{r_{t}}\!\left(\mathbf{I}-\hat{\mathbf{u}}_{t}\hat{\mathbf{u}}_{t}^{\top}\right)+\dfrac{1}{r_{r,i}}\!\left(\mathbf{I}-\hat{\mathbf{u}}_{r,i}\hat{\mathbf{u}}_{r,i}^{\top}\right)\right],&i\geq 1\\ \dfrac{2}{\lambda}\,\mathbf{v}^{\!\top}\!\left[\dfrac{1}{r_{r,0}}\!\left(\mathbf{I}-\hat{\mathbf{u}}_{r,0}\hat{\mathbf{u}}_{r,0}^{\top}\right)\right],&i=0\end{cases}\!.$
Finally, the FIM contribution from the $i$th radar reduces to $\mathbf{G}_{i}^{\top}\boldsymbol{\Sigma}_{i}^{-1}\mathbf{G}_{i},$ and the total effective FIM for the target position and velocity parameters is obtained by summing the individual contributions and the lower bound on the estimation error for parameters $\mathbf{p}$ and $\mathbf{v}$ is given by
$\operatorname{CRLB}(\begin{bmatrix}\mathbf{p}&\mathbf{v}\end{bmatrix}^{\top})=\mathbf{J}(\boldsymbol{\psi})^{-1}.$ (19)
@@ V-1 Frequency Synchronization
The leader broadcasts a TTW, where the tones are separated by a fixed reference frequency [1, 2, 15, 5]. Each follower extracts the reference clock signal using a dedicated circuit and uses it as a reference for clock synchronization. The TTW generated at the leader is also based on the same reference frequency. This ensures that the clocks and carriers at all radars tick at the same rate. However, clock offsets may still remain unaligned.
@@ VI Simulation Results
In this section, we evaluate the performance of the MSR system with and without follower radars. The simulation setup, illustrated in Fig. 2, consists of equally spaced one leader and 24 followers (along with follower index) each equipped with single antenna (resulting in total 25 antennas). A free-space path-loss model is assumed with a radar cross section (RCS) of $0.1~\text{m}^{2}$. The key radar parameters are summarized in Table I.
@@
Leader / Follower transmit power, $P^{L}_{\text{sync}}/P^{F}_{\text{sync}}$ 23/23 dBm
@@
For the leader–follower synchronization links, an Urban Macro (UMa) large-scale path-loss model with Rician fading is adopted [10], capturing both dominant LOS and diffuse NLOS components (exponential delay power profile). Depending on the node heights, the channels are categorized as Air-to-Ground (A2G) or Air-to-Air (A2A). As the radar altitude increases, the LOS component becomes stronger, leading to higher SNR and reduced multi-path effects. A Rician factor of 3 is used based on [10], with both leader and followers placed at a height of 10 m (A2A configuration). For frequency synchronization, the leader broadcasts a TTW with 10 MHz tone separation. Each follower extracts a reference clock from the received waveform and uses it to drive its internal circuitry. The frequency difference between the extracted reference and the nominal 10 MHz tone spacing is estimated and multiplied by the PLL (Phase-locked loop) factor to obtain the CFO error. The PLL factor refers to the multiplication ratio that a PLL applies to its reference input frequency to generate the carrier. For time synchronization, followers initiate a bidirectional waveform exchange at a transmit power of $P_{\text{sync}}^{F}$, and the leader responds accordingly. Since clock alignment depends on delay estimation (independent of the TOA estimation used for target tracking), the clock estimation error is modeled as twice the delay estimation error, accounting for estimation at both ends. An LFM waveform is used for delay estimation [17]. The CRLB for the first-path delay estimate is
@@
where $\zeta_{f}^{2}$ is the mean-squared bandwidth, the (pre-processed) $\mathrm{SNR}_{\mathrm{sync}}=P_{\mathrm{sync}}/\sigma_{n}^{2}$ with transmit power $P_{\mathrm{sync}}$ and noise variance $\sigma_{n}^{2}$, and $|h^{1}_{\mathrm{sync}}|^{2}$ is the dominant-path channel gain. For an LFM pulse of bandwidth $B_{\mathrm{sync}}$, $\zeta_{f}^{2}=\frac{(\pi B_{\mathrm{sync}})^{2}}{3}.$ Thus, the standard deviations of the frequency $\delta f$ and clock time $\delta t$ errors are computed and used as priors in the FIM calculation for delay and Doppler estimation at each radar. A time-bandwidth product of $B_{\mathrm{sync}}T_{\mathrm{sync}}=128$ is assumed.
The standard deviations of the CFO errors and clock error are shown in the Fig. 3. The maximum Doppler shift in the synchronization links is considered to be 0.1 Hz, arising from environmental disturbances. The performance plots show the clock error in nanoseconds (ns), with higher errors observed for followers farther from the leader and lower errors for those closer to it, indicating that higher SNR yields smaller clock errors. Increasing the $B_{\mathrm{sync}}$ further reduces the clock error. The CFO errors are primarily influenced by Doppler shifts and diffuse multi-path components in the environment. As multi-path severity increases, the likelihood of one or both tones experiencing deep fades also rises. Lower delay spreads correspond to higher coherence bandwidths. Hence, higher CFO errors are observed for delay spreads exceeding 100 ns.
@@
Fig. 4 compares the MSR performance under ideal conditions and with synchronization offsets introduced by the integrated synchronization technique. The PEB (position error bound) and VEB (velocity error bound) are obtained from (19), computed as the square root of the first three and the next three diagonal terms, respectively, with target locations and velocities uniformly drawn from the ranges in Table III. Performance improves with increasing $B_{\mathrm{sync}}$ used for delay estimation, while stronger diffuse multi-path (increasing delay spread $\tau_{rms}$) degrade accuracy. As synchronization-link performance improves, the MSR approaches the ideal case. The PEB is also compared with that of the monostatic radar [7] (equipped with antenna array of $3\times 3\times 3$ antenna array, resulting in total of 27 antenna) for fair comparison. The VEB remains unaffected by the synchronization bandwidth, as it is only influenced by the CFO errors resulting from frequency synchronization.